A fox fleeing from a hunter encounters a 0.785 m tall fence and attempts to jump it. The fox jumps with an initial velocity of 7.60 m/s at an angle of 45.0°, beginning the jump 1.96 m from the fence. By how much does the fox clear the fence? Treat the fox as a particle.
Answer #1

Initial speed has 2 components:
Vx=7.60*cos(45)
Vy=7.60*sin(45)

Horizontal motion is at constant speed of Vx, thus
d = Vx*t
t = d/Vx
t = 1.96/(7.60*cos(45))
t = 0365 s
vertical motion equation is (upward is positive)
h = -(1/2)*g*t^2 + Vy*t
h = -(1/2)*9.8*(0.365)^2 + (7.60*sin(45))*(0.365)
h = 1.31m

The fox will be
(1.31-0.785) m = 0.525 m ---> Answer
above the fence.

I hope help you !!